### v1

# Branch, or Layer? Zeroth-Order Optimization for Continual Learning of Vision-Language Models

Abstract
Continual learning in vision-language models (VLMs) faces critical challenges in balancing parameter efficiency, memory consumption, and optimization stability. While First-Order (FO) optimization (e.g., SGD) dominate current approaches, their deterministic gradients often trap models in suboptimal local minima and incur substantial memory overhead. This paper pioneers a systematic exploration of Zeroth-Order (ZO) optimization for vision-language continual learning (VLCL). We first identify the incompatibility of naive full-ZO adoption in VLCL due to modality-specific instability. To resolve this, we selectively applying ZO to either vision or language modalities while retaining FO in the complementary branch. Furthermore, we develop a layer-wise optimization paradigm that interleaves ZO and FO across network layers, capitalizing on the heterogeneous learning dynamics of shallow versus deep representations. A key theoretical insight reveals that ZO perturbations in vision branches exhibit higher variance than language counterparts, prompting a gradient sign normalization mechanism with modality-specific perturbation constraints. Extensive experiments on four benchmarks demonstrate that our method achieves state-of-the-art performance, reducing memory consumption by 89.1% compared to baselines. Code will be available upon publication.
1 Introduction
Continual Learning (CL) has witnessed significant advancements in convolutional architectures (e.g., ResNet erd ; res_2 ; res_3 ; pose and ViTvit_1 ; vit_2 ; vit_3 ). Recently, Vision-Language Models (VLMs)-based CL approaches have attracted growing research attention. Particularly, CLIP-based methods clip_cl_1 ; clip_cl_2 have demonstrated robust continual learning capabilities. For instance, ZSCL zscl introduces an unlabeled reference dataset for knowledge distillation to guide parameter space optimization. Mod-X mod-x maintains consistency in multimodal representation spaces through specially designed loss functions, effectively mitigating performance degradation caused by geometric variations in feature spaces. However, these methods require full-parameter fine-tuning of CLIP models, which incurs substantial computational overhead. The recently proposed MoE4Adapter MoE4Adapters addresses this limitation by adopting Parameter-Efficient Fine-Tuning (PEFT) strategies peft ; peft1 ; peft2 , achieving competitive CL performance with significantly reduced resource consumption.
Notably, existing VLM-based CL methods predominantly employ first-order (FO) optimization strategies (e.g., SGD, Adam) overview_opti . While these methods provide stable gradient descent directions, their inherent local convergence characteristics often lead to suboptimal solutions zo_intro_1 ; zo_intro_2 ; zo_intro_3 . MeZO mezo and MaZO maskzo both leverage ZO optimization to address the memory challenges of fine-tuning large language models in resource-constrained environments while maintaining or even surpassing the performance of traditional backpropagation-based fine-tuning. This breakthrough has inspired novel research directions in CL. For example, ZeroFlow zeroflow systematically compares various FO optimization strategies and discovers that ZO optimization exhibits unique advantages in alleviating catastrophic forgetting, providing new insights for optimizer design in continual learning systems.
However, the application of zero-order (ZO) optimization in Vision-Language Continual Learning (VLCL) remains understudied. Therefore, we conduct an in-depth exploration of ZO in VLCL using the state-of-the-art MoE4Adapter model MoE4Adapters as the baseline. First, we naively replace all FO optimizers with ZO counterparts to observe the direct impact of ZO in VLCL. However, experiments reveal that this approach causes excessive training divergence and failure to converge. Inspired by the cross-modal interaction consistency theory in VLMs cross_1 ; cross_2 , we introduce two investigations:
Investigation 1: Can ZO and FO collaborate for global model optimization across branches?
We implement ZO in either the visual or linguistic modality branch while retaining FO in the other. Experiments show that this strategy significantly outperforms the naive full-ZO approach.
Investigation 2: Can ZO and FO collaborate for hierarchical optimization within branch layers?
We thus deploy ZO only in specific layers (e.g., odd or consecutive layers) within a single branch while keeping FO in the remaining layers. Experiments further demonstrate that this approach yield better results, with certain configurations even surpassing the full-FO baseline. We attribute this to the stable directivity of FO optimization combined with the limited volatility of ZO, which helps the model escape local optima during optimization. Additionally, in layer-wise experiments, we find the visual branch exhibit significantly stronger fluctuations under ZO than the language branch. Therefore, we propose a stabilization mechanism into the visual branch to constrain ZO’s search range, further improving optimization convergence and overall continual learning performance.
In summary, the main contributions of this work are:
  • We reveal differential effects of ZO optimization on visual and language branches in VLCL and proposed a modality-specific framework combining ZO with FO optimization to balance stability and exploration.
  • Our layer-wise granular approach selectively deploys ZO in specific layers alongside FO in others, leveraging their synergy to escape local optima and enhance optimization efficiency.
  • Addressing ZO-induced volatility in the visual branch, we introduce a targeted stabilization mechanism to constrain its search range, significantly improving convergence stability and overall VLCL performance.
2 Optimization in Continual Learning: A Primer
Traditional optimization methods in continual learning primarily rely on FO gradient descent fo_1 ; fo_2 . Given a loss function $\mathcal{L}(\theta)$ and parameters $\theta$, FO updates parameters via:
$\theta=\theta-\eta\cdot\nabla_{\theta}\mathcal{L}(\theta),$ (1)
where $\nabla_{\theta}\mathcal{L}(\theta)$ is the FO gradient, and $\eta$ is the learning rate. While FO methods offer accurate gradient directions and have been widely adopted due to their efficiency, they may incur considerable memory overhead when higher-order derivatives (e.g., Hessians in second-order extensions) are involved lora ; memory_1 . Moreover, their deterministic update paths tend to limit exploration during training, potentially increasing the risk of convergence to suboptimal local minima and reducing adaptability in dynamic continual learning scenarios.
In contrast, ZO optimization estimates gradients through forward passes through purposeful perturbations. For a parameter subset $\theta^{k}_{m}$ (e.g., layer $k$ in modality $m$), ZO approximates gradients via directional perturbations:
$\nabla_{\text{ZO}}\mathcal{L}(\theta^{k}_{m})\approx\frac{\mathcal{L}(\theta^{k}_{m}+\varepsilon\Delta)-\mathcal{L}(\theta^{k}_{m})}{\varepsilon}\cdot\Delta,$ (2)
where $\Delta$ is a random directional vector, and $\varepsilon$ is a small perturbation scale. ZO methods avoid memory-intensive backward computations and introduce gradient stochasticity derivative , which may improve the ability to explore non-convex loss landscapes and escape poor local minima. However, the reliance on randomized perturbations can lead to variance in gradient estimates, whose effect may vary across different model architectures or modalities. These characteristics suggest that a hybrid optimization strategy, which balances FO’s directional precision with ZO’s exploratory potential, could offer a complementary pathway for improving training robustness in continual learning scenarios.
Figure 1: Illustration of our empirical study. The language encoder and vision encoder of CLIP are frozen, only the trainable units attached to each layer is performed to parameters update. To sum up, we systematically explores how ZO optimization operates in VLCL, including branches (Dual, Vision, or Language) and layers (w/ Hop-odd, w/ Hop-even, w/ Prefix (six) and w/ Suffix (six)).
3 Empirical Study of Zeroth-Order Optimization in VLCL
3.1 Implementation
Datasets and task construction. We evaluate our method on three datasets: CIFAR-100 (100 fine-grained classes), Tiny-ImageNet (200-class subset of ImageNet), and ImageNet-R (45,000 artistic variants spanning 200 ImageNet classes). For task construction under the CIL paradigm, we adopt the IncX configuration, where a dataset is split into sequential tasks with $X$ classes per task (e.g., Inc20 denotes 5 tasks with 20 classes each on CIFAR-100). This configuration is consistently applied across all datasets. All tasks enforce disjoint class distributions and exclude task-specific identifiers during inference to ensure a rigorous evaluation protocol.
Method. We build on the MoE4Adapter MoE4Adapters , a state-of-the-art VLCL framework that incorporates MoE modules into pre-trained vision-language models (e.g., CLIP). To explore PEFT alternatives, we introduce a variant, MoE4Adapter†, which replaces the original MoE modules with single-layer Low-Rank Adaptation (LoRA lora ) modules. This modification streamlines the architecture by reducing computational overhead while preserving adaptability, allowing direct comparison between MoE and LoRA within a unified framework.
Implementation details. All experiments employ the CLIP-ViT-B/16 backbone architecture, consistent with the baseline MoE4Adapter framework, to ensure direct comparability. Following MoE4Adapter’s paradigm, the pre-trained VLM backbone remains frozen, with only task-specific adapters (MoE or LoRA modules) being trainable. For the ZO-based method, we adopt a more conservative ZO strategy to enhance training stability and reduce the risk of noisy updates, which evaluates multiple candidate updates and selectively applies the one that yields the lowest loss. Hyperparameters including perturbation scale $\varepsilon=0.001$ for ZO gradients and FO/ZO mixing ratio $\lambda=1$, are validated on the first task and retained for subsequent tasks.
3.2 Rethinking Zeroth-Order Optimization in VLCL
The potential of zeroth-order optimization in VLCL. VLCL harnesses trainable units in VLMs to achieve parameter-efficient adaptation. Notably, trainable units often encounter optimization difficulties due to modality-specific feature distribution disparities. Conventional FO optimization method relies on precise gradient descent, hence leading to the attraction by local optima and suboptimal convergence. Besides, the necessity of back propagation requires storing high-gradients which significantly improve the memory overhead.
To mitigate these challenges, we propose integrating ZO optimization into VLCL. By leveraging its perturbation-based search mechanism, ZO enhances exploration in the parameter space, enabling escape from local optima. Furthermore, ZO eliminates the need for storing high-order gradient information, significantly reducing memory overhead—a critical advantage for resource-constrained CL tasks. However, does naively replacing FO with ZO necessarily lead to better performance?
Table 1: How ZO optimization affects VLCL through different branches (CLIP). w/ ZO denotes the branch (Dual, Vision, or Language) where ZO optimization is applied. †indicates that the expert modules in MoE4Adapter are replaced with LoRA.
Method CIFAR-100 Inc20 CIFAR-100 Inc10 Tiny-Imagenet Inc20 Image-R Inc20
Last. Avg. Last. Avg. Last. Avg. Last. Avg.
MoE4Adapter MoE4Adapters 80.47 86.97 77.52 85.21 52.13 60.55 65.36 71.53
      Dual w/ ZO 69.29 77.36 67.64 75.88 42.40 47.64 58.56 65.92
      Vision w/ ZO 76.05 83.93 72.98 82.08 49.65 57.90 62.54 69.84
      Language w/ ZO 80.94 87.00 76.74 85.03 49.14 58.69 64.38 70.38
MoE4Adapter† 80.44 87.10 79.66 86.34 51.93 59.80 64.34 71.79
      Dual w/ ZO 71.07 78.37 69.70 76.82 44.53 52.71 58.71 65.97
      Vision w/ ZO 75.86 83.86 73.73 82.25 49.98 58.11 62.70 69.88
      Language w/ ZO 77.47 85.40 79.63 87.01 49.90 58.92 64.33 70.59
Empirical analysis of zeroth-order optimization failure in VLCL. We attempt to apply ZO into the VLCL, a straightforward intuition is to replace FO optimizers with ZO methods across both the vision and language branches including all trainable units, and the results are shown in Table 1. However, it can be observed when ZO is adopted in both vision and language branch (dual w/ ZO), the performances are significantly degraded regardless of MoE or LoRA settings, with “Last.” decreasing by approximately 6%–11% and "Avg." by approximately 6%–13%. It can be attributed to the reason that the variance of ZO destabilizes VLCL training, leading to the optimization process is difficult to converge. To verify this point, we plot the loss function trajectories across these experimental settings shown in Figure 2. The Figure 2(a) and 2(b) reveal that dual-branch ZO approach suffers from severe loss oscillations and failed convergence compared to the original FO optimization. From this observation, we conclude that full ZO optimization is fundamentally ill-suited for VLCL, as its inherent gradient estimation fluctuations induce training instability. On the contrary, FO optimization provides stable gradient directions throughout the training process. This raises the question: Can synergistic integration of ZO and FO optimization achieve better performance?
(a) Dual w/ FO
(b) Dual w/ ZO
(c) Vision w/ ZO
(d) Language w/ ZO
Figure 2: How ZO optimization affects loss convergence of VLCL across different branches (CLIP). w/ ZO denotes the branch (Dual, Vision, or Language) where ZO optimization is applied.
How can zeroth-order optimization be effective in VLCL? Branches, or Layers? We further investigate a heterogeneous optimization strategy: applying ZO to only one modality branch while retaining FO for the other. From a qualitative perspective, Figure 2(c) and 2(d) reveal when ZO is adopted into a single branch (vision or language), the overall loss function trajectory is promising to converge compared with dual branch ZO. We consider that the single-branch FO provides optimization stability, guiding the overall training process maintains consistency. Meanwhile, we can find that the performance obtains significant improvement shown in Table 1, most of the results are close to the baseline, and even some results outperform baseline. It can be explained that ZO’s perturbation-based gradient estimation introduces controlled stochasticity into the optimization process, which probabilistically assists the optimizer in evading suboptimal local minima. Additionally, we find that the performance of language w/ ZO generally outperform vision w/ ZO, we argue that the optimization stability of language branch is stronger than vision branch, which is verified by the loss trajectories of Figure 2(c) and 2(d). These findings suggest a ZO-FO synergy strategy which balances performance exploration and training convergence in VLCL, validating the feasibility of the integration optimization.
Building on these observations, the effectiveness of coarse-grained integration motivates us to explore the potential of synergistic optimization between ZO and FO methods. To further investigate this synergy, we consider a fine-grained perspective: Can layer-wise allocation of ZO and FO optimization within modality-specific enhance CL performance?
3.3 Why Layers Matter: Triggering Effective Continual Adaptation
To explore the performance regarding layer-wise allocation of ZO and FO, four layer-wise ZO patterns were tested: Hop-odd (adopt ZO in odd layers), Hop-even (adopt ZO in even layers), Prefix (six) (adopt ZO in first six layers) and Suffix (six) (adopt ZO in last six layers) in dual or single modality branch, with FO used in other layers. The selection of modality branch is still refer to Table 1.
(a) Dual w/ Hop-odd.
(b) Vision w/ Hop-odd.
(c) Language w/ Hop-odd.
Figure 3: Analyzing convergence behavior of VLCL in Hop-odd across Dual, Vision, Language.
Table 2: How ZO optimization affects VLCL through different layers (CLIP). We design four configurations across layers from different branches: w/ Hop-odd (ZO in odd layers), w/ Hop-even (ZO in even layers), w/ Prefix (six) (ZO in first six layers) and w/ Suffix (six) (ZO in last six layers).
Method CIFAR-100 Inc20 CIFAR-100 Inc10 Tiny-Imagenet Inc20 Image-R Inc20
Last. Avg. Last. Avg. Last. Avg. Last. Avg.
MoE4Adapter MoE4Adapters Dual w/ ZO 69.29 77.36 67.64 75.88 42.40 47.64 58.56 65.92
      w/ Hop-odd 81.64 88.11 79.22 86.98 51.68 59.75 65.24 71.59
      w/ Hop-even 81.59 88.07 79.12 86.81 51.95 60.53 64.99 71.29
      w/ Prefix (six) 79.12 85.89 76.53 84.77 50.47 59.34 65.01 71.36
      w/ Suffix (six) 80.98 87.52 78.03 86.63 51.48 59.33 63.10 69.99
MoE4Adapter MoE4Adapters Vision w/ ZO 76.05 83.93 72.98 82.08 49.65 57.90 62.54 69.84
      w/ Hop-odd 81.83 88.36 79.39 86.96 51.56 60.11 65.68 72.05
      w/ Hop-even 82.41 88.36 78.95 86.74 52.27 60.98 64.99 72.10
      w/ Prefix (six) 79.34 86.17 76.72 84.85 50.14 59.73 65.25 71.96
      w/ Suffix (six) 82.21 88.33 79.23 87.03 51.65 60.20 64.25 70.76
MoE4Adapter MoE4Adapters Language w/ ZO 80.94 87.00 76.74 85.03 49.14 58.69 64.38 70.38
      w/ Hop-odd 82.28 88.51 79.28 87.05 51.73 60.59 65.20 71.90
      w/ Hop-even 82.17 88.45 78.73 86.27 52.07 60.71 65.06 72.16
      w/ Prefix (six) 82.19 88.51 79.07 86.83 52.02 60.49 64.87 71.64
      w/ Suffix (six) 82.19 88.30 78.82 86.71 51.57 60.38 65.17 72.10
Layer-wise ZO unlocks VLCL performance potential. We then apply layer-wise ZO optimization to both dual and single modality branches, with the remaining layers optimized by FO. As shown in Table 2, we observe that layer-wise strategy can provide significant performance improvements compared to applying ZO across all trainable units in branches. In dual w/ ZO, the layer-wise ZO optimization averagely improves 9.4% accuracy on four patterns across all dataset. More interestingly, we observe that certain fine-grained layer-wise ZO patterns can outperform full FO approaches. To investigate the layer-wise effectiveness, we first analyze the performance of applying the collaborative ZO-FO optimization to a uniform layering strategy (with Hop-odd selected as a representative case). The loss trajectories under different modalities are recorded and presented in Figure 3, it demonstrates a more stable convergence process during training compared to that shown in Figure 2. We argue that more fine-grained layer-wise strategy further amplifies the respective advantages of ZO and FO in VLCL, FO provides stable gradient directions and ZO’s stochastic perturbations help the optimization process escape local minima.
Observation on Layer-wise Heterogeneity in ZO for VLCL. To gain deeper insights, we start to analyze the impact of different layer-wise settings on performance. Interestingly, we observe that all SOTA results emerge when ZO and FO optimization are applied in an interleaved manner across layers (e.g., on Hop-odd or Hop-even). To further investigate this phenomenon, we analyze the gradient behavior under four different layer-wise configurations within the language branch. The corresponding gradient distributions are visualized in Figure 4. We clearly observe that the gradient variance is significantly lower when ZO and FO are interleaved across layers, compared to configurations where either ZO or FO is applied continuously throughout. We hypothesize that this benefit stems from the functional heterogeneity across layers: shallow layers focus on local features, while deeper layers capture abstract semantics. A uniform optimization method may overlook such diversity, whereas interleaving ZO and FO better aligns with each layer’s exploration and stability needs, leading to a more robust optimization process that stabilizes gradient flow while facilitating escape from local minima.
(a) Prefix (six).
(b) Suffix (six).
(c) Hop-odd.
(d) Hop-even.
Figure 4: How ZO optimization affects gradient variance across layers in VLCL.
(a) Dual w/ FO
(b) Dual w/ ZO
(c) Vision w/ ZO
(d) Language ZO
Figure 5: Analyzing gradient variance of VLCL in Hop-odd across Dual, Vision, Language.
(a) Dual w/ FO
(b) Dual w/ ZO
(c) Vision w/ ZO
(d) Language w/ ZO
Figure 6: Analyzing gradient variance of VLCL in Hop-even across Dual, Vision, Language.
3.4 A New Enhancement from Vision Branch Discrepancy
Understanding the discrepant behavior of ZO in vision branch. From Figure 2, we observe that when ZO is adopted in a single branch, the loss convergence trajectory of visual branch demonstrates significantly inferior performance compared with language branch. To further investigate this phenomenon, we record the gradient variance distribution of different optimization strategies from a layer-wise perspective in VLCL, since layer-wise ZO obtain better performance, and the visualization results are shown in Figure 5 and 6. It can be observed that dual branch FO optimization exhibit a minimal numerical fluctuations in gradient variance, reflecting robust convergence stability regardless of Hop-odd or Hop-even layers setting, while the lack of sufficient fluctuation might lead the optimization process to converge to local optima. In contrast, dual branch ZO induce severe oscillations, causing the optimization trajectory to diverge. When the ZO-FO collaborative mechanism is employed, we interestedly observe that under the same layer wise setting, the gradient variance of the visual branch ZO is also more violent than that of language. This observation motivates a critical inquiry: Could targeted suppression of ZO-induced perturbations in the visual modality yield performance gains?
Table 3: Effect of vision branch discrepancy on performance. Dual w/ Hop-odd/even indicates the results of adopting Hop-odd and Hop-even layer-wise ZO in dual branch.
Method CIFAR-100 Inc10 Tiny-ImageNet Inc20 Image-R Inc20
Last. Avg. Last. Avg. Last. Avg.
Dual w/ Hop-odd 79.22 86.98 51.68 59.75 65.24 71.59
Vision Discrepancy 79.36 87.02 52.35 59.75 65.32 71.93
Dual w/ Hop-even 79.12 86.81 51.95 60.53 64.99 71.29
Vision Discrepancy 79.87 87.25 52.46 61.23 65.80 71.82
Gradient Regularization and Vision Branch Perturbation Control. To validate this hypothesis, we propose a ZO optimization strategy that incorporates gradient regularization during the estimation process, with explicit constraints on perturbations applied to the vision branch to mitigate instability. Specifically, we first introduce a signed gradient transformation to unify the gradient values which can be expressed as:
$\tilde{g}=\operatorname{sign}(g),\quad\text{where}\quad\operatorname{sign}(g_{i})=\begin{cases}+1,&\text{if }g_{i}>0\\ 0,&\text{if }g_{i}=0\\ -1,&\text{if }g_{i}<0\end{cases}\quad\forall i\in\{1,\dots,d\},$ (3)
where $g$ denotes the original gradient, and $\tilde{g}$ is the transformed signed gradient. This transformation retains only the direction information of the gradient and discarding the amplitude information. Furthermore, we implement modality-specific perturbation factors for ZO optimization, assigning a deliberately lower value to the vision branch compared to the language branch. This strategy theoretically reduces the instability of ZO optimization and suppresses gradient fluctuations in the vision modality. To verify this strategy, we first record the gradient variance on CIFAR-100, and plot the comparison shown in Figure 7. It can be observed that our strategy significant reduce the vision branch gradient fluctuations expressed in Figure 7(b) and 7(d) (shown in red line), regardless of Hop-odd or Hop-even. We further validate this conjecture through quantitative analysis shown in Table 3, it can be find that stabilizing ZO gradient estimation and reducing vision branch perturbations yields consistent performance gains across CIFAR-100, Tiny-Imagenet and Image-R datasets. This method focus on the instability limits of ZO in VLCL and maintains the balance of the optimization process across different modality branches, paving the way to further refines the application strategy of ZO optimization in VLCL.
(a) Hop-odd
(b) Hop-odd
(c) Hop-even
(d) Hop-even
Figure 7: Analyzing the effect of vision discrepancy on gradient variance in Hop-odd vs. Hop-even. Blue shows original results, while red shows the positive impact attributed to vision discrepancy.
4 Ablation, Empirical Demonstration and Beyond
Exploring diverse ZO strategies. Our ablation start to explore other ZO strategies of hybrid ZO-FO collaboration in VLCL. As shown in Table 4, applying ZO to dual branches significantly degrades performance compared to the baseline, confirming the instability caused by excessive gradient oscillations. In contrast, single-branch ZO optimization mitigates this issue, with the significant improvement in Last. and Avg.. We then explore the impact of different ZO optimization on VLCL performance. As we default to the conservative ZO strategy in the main analysis, we further conduct more aggressive ZO variants. Specifically, we examine a naive ZO approach—referred to as ZO* which performs a single gradient estimation and directly updates the parameters without any loss-based validation. On top of this, we introduce a variant named Sign, which incorporates a signed gradient transformation to regulate the magnitude of the estimated gradients. This design aims to mitigate the instability caused by ZO estimates while preserving directional information.
Table 4: Behavioral consistency across diverse ZO strategies. * indicates a naive ZO strategy. And Sign indicates a ZO strategy using the gradient transformation method.
Method Strategy CIFAR-100 Inc10
Last. Avg.
MoE4Adapter
      Dual w/ ZO ZO* 66.67 75.08
      Vision w/ ZO ZO* 72.32 81.64
      Language w/ ZO ZO* 76.19 84.58
      Dual w/ ZO Sign 66.68 77.20
      Vision w/ ZO Sign 73.34 83.47
      Language w/ ZO Sign 78.52 85.91
It can be observed that the aggressive ZO* optimization degrades performance, as it relies on a single gradient estimation without validating the update direction, making it more sensitive to the fluctuations in the loss landscape compared with conservative ZO. Notably, adopting Sign-based gradient estimation enhances performance across all configurations, suggesting that enforcing gradient amplitude consistency in ZO perturbations helps stabilize the optimization trajectory and improves convergence. This observation aligns with our proposed ZO optimization strategy in Section 3.4, which advocates for gradient magnitude control as a means to suppress fluctuation updates and maintain optimization balance across modalities.
(a) Last. of CIFAR-100.
(b) Avg. of CIFAR-100.
(c) Last. of Image-R.
(d) Avg. of Image-R.
Figure 8: Significance analysis of performance across Dual, Vision, Language.
Significance analysis. We then explore the average incremental performance among 5 runs on CIFAR-100 and Image-R, with 3 different ZO configurations (dual branch, vision branch and language branch), recording the accuracy of Last. and Avg. shown in Figure 8. We observe that ZO optimization yields smaller performance variance on the CIFAR-100 dataset than on Image-R across different configurations, regardless of Last and Avg. One possible explanation is that Image-R derived from the larger-scale ImageNet, contains more fine-grained and semantically diverse categories than CIFAR-100. This increased label complexity leads to a more rugged and high-dimensional loss landscape,
Table 5: Comparison of GPU memory usage across branches(CLIP).
Method Memory $\downarrow$
MoE4Adapter $\sim$19.96GB
      Dual w/ ZO $\sim$2.17GB
      Vision w/ ZO $\sim$6.93GB
      Language w/ ZO $\sim$12.39GB
MoE4Adapter† $\sim$15.11GB
      Dual w/ ZO $\sim$1.73GB
      Vision w/ ZO $\sim$5.71GB
      Language w/ ZO $\sim$11.09GB
which makes ZO gradient estimation on stochastic perturbations harder to stabilize. Additionally, from the perspective of different ZO configurations, applying ZO optimization to the language branch consistently yields the better performance. We hypothesize that this is because the language branch typically operates on lower-dimensional tensors, making it less susceptible to the instability introduced by random perturbations in ZO gradient estimation. In contrast, the higher-dimensional vision branch benefits more from FO gradients, which provide precise gradient optimization direction. The combination of FO-dominated updates in the vision branch and ZO exploration in the language branch may facilitate better escape from local minima and lead to more robust convergence.
Quantifying memory efficiency. We record the memory usage across different ZO settings, as shown in Table 5. Compared to applying FO optimization in the dual branch MoE4Adapter, using ZO optimization leads to a 89.1% reduction in memory consumption, due to the elimination of gradient backpropagation and storage overhead. When applying ZO optimization to a single branch, memory consumption is reduced by 65.3% on the visual branch and 37.9% on the language branch, respectively. The greater memory reduction on the visual branch is attributed to its higher input dimensionality, as it processes image data. Applying ZO in this context helps to alleviate memory pressure by avoiding storage of high-dimensional gradients. For the results in LoRA architecture, it inherently requires fewer trainable units to perform parameters update, resulting in overall lower memory consumption. Overall, these findings highlight ZO’s inherent advantage in significantly reducing memory consumption, especially in high-dimensional settings, making it a highly practical and efficient optimization alternative for resource-constrained VLCL scenarios where memory efficiency is critical.
5 Related Work
Continual learning of VLM. Recent advances in CL have witnessed the emergence of VLMs as promising solutions to mitigate catastrophic forgetting through their generalized multimodal representations. Pioneering studies lwf-vr ; clip_cl2 demonstrate that pretrained VLMs like CLIP clip inherently possess remarkable continual learning capabilities even without fine-tuning. PROOF vlm_lwf further enhances CL robustness by integrating multimodal cues with adaptive mapping strategies. Nevertheless, conventional VLM-based approaches predominantly emphasize task-specific feature acquisition for new domains ralf , inadvertently compromising the integrity of previously learned representations and leading to progressive performance degradation. To address this limitation, recent efforts adopt PEFT techniques peft1 ; peft2 that selectively update lightweight modules while maintaining frozen backbone parameters. MoE4Adapters MoE4Adapters introduces a mixture-of-experts (MoE) architecture, enabling task-specific feature specialization without cross-task interference and achieving state-of-the-art performance. However, these methods universally rely on FO optimization, inherently restricting their capacity to explore optimal parameter trajectories during optimization.
Optimization for continual learning. From an optimization view, existing CL methods predominantly focus on reconciling the stability-plasticity dilemma staplast through gradient technology unigrad . Orthogonal parameter updates grad1 ; grad2 ; grad3 ; grad4 and sharpness-aware minimization flat1 ; flat2 ; flat3 represent two mainstream directions, which respectively aim to decouple task-specific gradients and converge to flat loss minima for improved generalization. While these strategies enhance optimization stability, they often inadequately address the exploration-exploitation trade-off, as deterministic gradient descent trajectories tend to converge to suboptimal local minima with limited perturbation resilience. Emerging ZO optimizationzo1 ; zo2 ; zeroflow presents a paradigm shift by employing gradient-free stochastic perturbations to estimate descent directions. This approach offers dual advantages: 1) inherent stochasticity facilitates escape from local optima through controlled parameter space exploration, and 2) elimination of backward propagation reduces memory overhead by avoiding gradient matrix storage.
Our work. This work systematically investigates how ZO optimization can be effectively adapted to VLCL, addressing catastrophic forgetting by harmonizing the inter-modal asymmetry of VLMs with the stochastic perturbations of ZO optimization.
6 Conclusion
In this work, we systematically investigate ZO optimization into VLCL and propose a novel hybrid ZO-FO optimization paradigm. Through extensive empirical analysis, we identify two critical challenges of applying ZO in VLCL: destabilized training caused by excessive gradient variance and modality-specific optimization discrepancies. To address these, we first demonstrate that selectively applying ZO to specific branches (vision or language) while retaining FO optimization in others significantly outperforms naive full-ZO approaches. Building on this, we further propose a layer-wise collaborative strategy that interleaves ZO and FO across network layers, achieving state-of-the-art performance by harmonizing stochastic exploration with deterministic refinement.
Limitation. This work currently focuses on CLIP-based architectures and image-text modalities. The generalization of ZO-FO collaboration to other VLMs (e.g., multimodal transformers with audio or video inputs) remains unexplored, particularly in scenarios where modalities exhibit heterogeneous feature distributions or temporal dependencies.
Supplementary Material
In the main paper, we analyze the application of ZO optimization in VLCL, covering both branch-wise and layer-wise settings. Extensive evaluations are conducted based on the MoE4Adapter baseline, and further preliminary explorations are carried out under the LoRA-based setting. In the supplementary, we provide additional results to examine the consistency of our observations under the LoRA setting. This material is divided into two sections:
  • Section A investigates the performance of layer-wise allocation of ZO and FO optimization under the LoRA setting, building upon the performance trends and layer heterogeneity analyses presented in the main paper.
  • Section B explores the vision branch discrepancy of ZO optimization and evaluates the proposed method from the main paper within the LoRA setting.
Appendix A Investigating Layer-wise ZO optimization for VLCL under LoRA Setting
We give the performance of layer-wise ZO optimization with both dual and single modality branches under LoRA setting shown in Table 6. Specifically, we choose two layer configurations: Hop-odd and Suffix (six). It can be observed that the performance trends are consistent with those reported in the main paper, demonstrating the layer-wise ZO optimization remains effective under the LoRA setting. We also show the convergence behavior of Hop-odd on Image-R dataset in Figure 9, and find that the training process remains consistently stable throughout.
Table 6: How ZO optimization affects VLCL through different layers (CLIP) under LoRA based setting (represented by †). We choose two configurations across layers from different branches: w/ Hop-odd and w/ Suffix (six).
Method CIFAR-100 Inc20 CIFAR-100 Inc10 Tiny-Imagenet Inc20 Image-R Inc20
Last. Avg. Last. Avg. Last. Avg. Last. Avg.
MoE4Adapter†MoE4Adapters Dual w/ ZO 71.07 78.37 69.70 76.82 44.53 52.71 58.71 65.97
      w/ Hop-odd 79.36 86.93 78.76 86.14 52.04 60.47 64.90 71.62
      w/ Suffix (six) 79.38 86.55 77.88 85.67 51.56 60.17 63.15 70.34
MoE4Adapter†MoE4Adapters Vision w/ ZO 75.86 83.86 73.73 82.25 49.98 58.11 62.70 69.88
      w/ Hop-odd 79.92 87.13 79.16 86.53 51.94 60.35 65.16 71.80
      w/ Suffix (six) 79.60 86.79 78.96 86.35 52.27 60.75 64.34 70.99
MoE4Adapter†MoE4Adapters Language w/ ZO 77.47 85.40 79.63 87.01 49.90 58.92 64.33 70.59
      w/ Hop-odd 79.82 86.90 79.88 87.38 52.32 61.06 65.59 71.70
      w/ Suffix (six) 79.68 86.94 78.46 86.13 52.22 60.61 65.51 71.67
(a) Dual w/ Hop-odd †.
(b) Vision w/ Hop-odd †.
(c) Language w/ Hop-odd †.
Figure 9: Analyzing convergence behavior of VLCL in Hop-odd across Dual, Vision, Language under LoRA setting (represented by †).
Additionally, we similarly observe that interleaved layer-wise ZO manner (Hop-odd) yields superior performance compared with consecutive layers under the LoRA setting. To better understand the results, we separately record the gradient distributions of the Hop-odd and Hop-suf layer-wise strategies across the dual, vision, and language branches on the Image-R dataset. Figure 10 clearly reveals that interleaved layer-wise ZO optimization yields more stable gradient behavior during training, suggesting interleaving of ZO and FO better meet each layer’s needs for gradient exploration and stability, which aligns with the hypothesis proposed in the main paper.
(a) Dual w/ Hop-suf †.
(b) Vision w/ Hop-suf †.
(c) Language w/ Hop-suf †.
(d) Dual w/ Hop-odd †.
(e) Vision w/ Hop-odd †.
(f) Language w/ Hop-odd †.
Figure 10: How ZO optimization affects gradient variance across layers under LoRA setting (represented by †).
Appendix B Exploring ZO Vision Branch Discrepancy under LoRA Setting
In the main paper, we explore the discrepant behavior of ZO across different modality branches, which motivates our proposed improvements for applying ZO in the VLCL setting. We replace the baseline architecture with LoRA and conduct the same investigation under identical experimental settings to revisit this phenomenon. Table 7 demonstrates that applying gradient regularization and reducing the perturbation in the vision branch leads to further improvements in most experimental results, this indicates that our proposed method remains effective under the LoRA setting. We further analyze this method by comparing the gradient variance in Hop-odd vs. Hop-even on image-R dataset shown in Figure 11, it can be observed that the proposed method has a positive impact, as the gradient variance becomes more stable. These results are broadly consistent with the conclusions presented in the main paper, further demonstrating the effectiveness of the proposed method.
Table 7: Effect of vision branch discrepancy on performance under LoRA setting (represented by †). Dual w/ Hop-odd/even indicates the results of adopting Hop-odd and Hop-even layer-wise ZO in dual branch.
Method CIFAR-100 Inc10 Tiny-ImageNet Inc20 Image-R Inc20
Last. Avg. Last. Avg. Last. Avg.
Dual w/ Hop-odd 78.76 86.14 52.04 60.47 64.90 71.62
Vision Discrepancy 78.84 86.11 52.12 60.67 64.92 71.67
Dual w/ Hop-even 79.22 86.27 52.66 61.32 65.68 71.67
Vision Discrepancy 79.30 86.23 53.54 61.99 65.89 71.98
(a) Hop-odd
(b) Hop-odd
(c) Hop-even
(d) Hop-even
Figure 11: Analyzing the effect of vision discrepancy on gradient variance in Hop-odd vs. Hop-even for Lora setting (represented by †). Blue shows original results, while red shows the positive impact attributed to vision discrepancy.